OpenStudy (anonymous):

solving linear equations 2/3x-18=x/6

OpenStudy (anonymous):

\[\large \frac{2}{3}x-18=\frac{x}{6}\]?

OpenStudy (anonymous):

i hate fractions multiply both sides by 6 and get \[4x-108=x\] then is should be much easier